OpenStudy (anonymous):

An investor invested a total of $1500 in two mutual funds. One fund earned a 4% profit value while the other earned 2% profit. If the investor's profit was $36, how much was in each mutual fund? The amount invested in the fund that earned 4% was __________. The amount invested in the fund that earned 2% was ________>

OpenStudy (jdoe0001):

let's say, the 2 amounts for each fund are called "a" and "b" the total amount he invested was $1500, that is, amount "a" plus amount "b" = 1500, that is \(\bf a + b = 1500\) the earned profits are, say 4% from "a", and 2% from "b", is $36, that is \(\bf \textit{to express the percentages in decimals}\\ 4\% = \cfrac{ 4 }{ 100 } \implies 0.04\\ 2\% = \cfrac{ 2 }{ 100 } \implies 0.02\) \(\bf 0.04\times a + 0.02\times b = 36 \implies 0.04 a + 0.02 b = 36\)

OpenStudy (jdoe0001):

if we know that say, a + b = 1500 , can you solve that for say... "b", what would it look like?

OpenStudy (anonymous):

I am lost that why I am asking for help its been over 13 years since i have done this

OpenStudy (jdoe0001):

well... how about just solving for "b" in a + b = 1500

OpenStudy (anonymous):

still dont get the problem

OpenStudy (jdoe0001):

http://www.youtube.com/watch?v=IIDTlgvo0DM